the measure of an angle is 30 degree more than twice the measure of its supplement. Find the measures of the angles.

Answer by psbhowmick(529)

Let the angle be X degree.
Then its supplement is (180-X) degree.
Given that, Angle - 2 x Supplement = 30.
X - 2(180 - X) = 30
X - 360 + 2X = 30
3X = 30 + 360 = 390
X = 390/3 = 130
Hence, the angles are 130 and 50 degrees.